70% of Valorites can answer the previous question correctly, suppose you select 3 Valorites at random, what is the probability that at least one of the 3 can answer correctly?
Answer by stanbon(75887)

70% of Valorites can answer the previous question correctly, suppose you select 3 Valorites at random, what is the probability that at least one of the 3 can answer correctly?
------
Binomial Problem with n = 3 and p(correct) = 0.7
----
P(1 <= x <= 3)
= 1 - P(x=0)
= 1- 3C0(0.7)^0(0.3)^(3)
= 1 - 0.3^3
= 0.973
=================================
Cheers,
Stan H.
=================================